OpenStudy (anonymous):

solving with variables on both sides: 8r = -4r-12 2y = 2+4y 3y+2y = 3y+6 4k-6k = 6-k

OpenStudy (anonymous):

1) 8r=-4r-12 8r+4r=-12 12r=-12 r=-12/12 = -1 r=-1 2)2y=2+4y 2y-4y=2 -2y=2 y=-1 3)3y+2y = 3y+6 5y-3y=6 2y=6 y=6/2=3 y=3 4)4k-6k = 6-k -2k=6-k -6=-k+2k -6=k